### v1

@@ Abstract
The evolution of AI and high-throughput technologies has boosted a rapid increase in the number of new materials, challenging our computational ability to comprehensively analyze their properties. Relaxed crystal structures often serve as the foundational basis for further property calculations. However, determining equilibrium structures traditionally involves computationally expensive iterative calculations. Here, we develop DeepRelax, an efficient deep generative model designed for rapid structural relaxation without any iterative process. DeepRelax learns the equilibrium structural distribution, enabling it to predict relaxed structures directly from their unrelaxed counterparts. The ability to perform structural relaxation in just a few hundred milliseconds per structure, combined with the scalability of parallel processing, makes DeepRelax particularly useful for large-scale virtual screening. To demonstrate the universality of DeepRelax, we benchmark it against three different databases of X-Mn-O oxides, Materials Project, and Computational 2D Materials Database with various types of materials. In these tests, DeepRelax exhibits both high accuracy and efficiency in structural relaxation, as further validated by DFT calculations. Finally, we integrate DeepRelax with an implementation of uncertainty quantification, enhancing its reliability and trustworthiness in material discovery. This work provides an efficient and trustworthy method to significantly accelerate large-scale computations, offering substantial advancements in the field of computational materials science.
@@ 1 Introduction
The discovery of new materials has significantly accelerated in the past decade, owing in part to the expansion of crystal structure databases by high-throughput density functional theory (DFT) calculationsong2013python ; saal2013materials ; curtarolo2012aflow ; zhou20192dmatpedia and advanced machine learning (ML) algorithms chen2021phase ; xie2021crystal ; zeni2023mattergen ; zhao2023physics . For example, very recently, around 2.2 million new crystal materials have been discovered by Google DeepMind using graph neural networksmerchant2023scaling .
@@
Structural relaxation, the first step and foundation for further computationally analyzing properties of any new materials, is typically achieved using iterative optimization, such as ab initio methods or ML algorithms as shown in Figure 1 (a). For example, each iteration in DFT calculations involves solving the Schrödinger equation to determine the electronic density distribution, from which the total energy of the system can be calculated. The forces on each atom, derived from differentiating this energy with respect to atomic positions, guide atomic movements to lower the system’s energy, typically using optimization algorithms. Despite its effectiveness, the high computational demands and poor scalability of DFT limit its widespread application across high-dimensional chemical and structural spaceszuo2021accelerating , such as the structural relaxation of the aforementioned 2.2 million new crystals or complex chemical reaction surfaces or doped semiconductor interfaces.
@@
On the other hand, machine learning (ML) has emerged as a promising alternative for predicting relaxed structureschen2022universal ; deng2023chgnet ; zuo2021accelerating ; mosquera2024machine ; kolluru2022open ; kim2023structure ; yoon2020differentiable ; wang2024concurrent ; Omee2024 . As conventional iterative optimization, iterative ML approacheschen2022universal ; deng2023chgnet ; zuo2021accelerating ; kolluru2022open ; mosquera2024machine ; wang2024concurrent ; Omee2024 utilize surrogate ML models to approximate energy and forces at each iteration, as shown in Figure 1 (a), thereby circumventing the need to solve the computationally intensive Schrödinger equation. These ML models can retain DFT-level accuracy by training on extensive databases containing detailed information on structural relaxations, including energy, forces, and stress.
@@
However, there are two primary challenges in current iterative ML structural optimizers: training data limitations and non-scalability. Their training dataset must include full or partial intermediate steps of DFT relaxation. However, almost all publicized material databases, such as ICSDbelsky2002new and 2DMatPediazhou20192dmatpedia , do not provide such structural information, potentially limiting the application of iterative ML methods. The other challenge is that the large-scale parallel processing capability of iterative ML methods is limited due to their step-by-step nature. To address this, Yoonyoon2020differentiable and Kimkim2023structure conceptually proposed direct ML approaches to predict the final relaxed structures directly from their unrelaxed versions. However, these approaches have only been validated on specific datasets or systems, and their universal applicability to diverse datasets or systems remains unproven.
@@
Figure 1: An overview of ML methods for crystal structure relaxation. Panel (a) depicts iterative ML methods that iteratively estimate energy and force to determine the equilibrium structure. Panel (b) illustrates our proposed DeepRelax method, which directly predicts the equilibrium structure without the iterative process.
@@
In this work, we develop a scalable, universal, and trustworthy deep generative model for fast structural relaxation, DeepRelax, by learning from the structural relationships between unrelaxed and relaxed crystals. DeepRelax requires only the initial crystal structures to predict equilibrium structures in just a few hundred milliseconds on a single GPU. Furthermore, DeepRelax can efficiently handle multiple crystal structures in parallel by organizing them into mini-batches for simultaneous processing. This capability is especially advantageous in large-scale virtual screening, where rapid assessment of numerous unknown crystal configurations is essential. Most importantly, DeepRelax employs an uncertainty quantification method to assess the reliability of the predicted structure, enabling the identification and exclusion of predictions that are likely to be unreliable. We evaluate DeepRelax across three different datasets, including diverse 3D and 2D materials: the Materials Project (MP)jain2013commentary , X-Mn-O oxideskim2020generative ; kim2023structure , and the Computational 2D Materials Database (C2DB)haastrup2018computational ; gjerding2021recent ; lyngby2022data . DeepRelax not only demonstrates superior performance compared to other direct ML methods but also exhibits competitive accuracy to the leading iterative ML model, M3GNetchen2022universal , while being approximately 100 times faster in terms of speed. Moreover, we conduct DFT calculations to assess the energy of DeepRelax’s predicted structures, confirming our model’s ability to identify energetically favorable configurations. Finally, we would like to highlight that the aim of using DeepRelax is not to replace DFT relaxation, but to make the predicted structures close to the final relaxed configuration. Thus, the DFT method can rapidly complete the relaxation with only a few steps, significantly speeding up the traditional ab initio relaxation process, especially for complex structures.
@@ 2.1 DeepRelax architecture
DeepRelax emerges as a solution to the computational bottlenecks faced in DFT methods for crystal structure relaxation. The foundation of DeepRelax is a periodicity-aware equivariant GNN (PaEGNN). This network is designed for handling periodic boundary conditions (PBCs) explicitly and ensures equivariance to actively explore crystal symmetries, thereby offering a more enriched geometrical representation for crystal structures. Figure 1 (b) shows the workflow of DeepRelax, which takes a crystal structure as input and uses PaEGNN to predict the relaxation quantities. These quantities include interatomic distances in the relaxed structure and their associated local uncertainties (i.e., bond-level uncertainties), displacements between the initial and relaxed structures (also quantified with local uncertainties), and lattice parameters of the relaxed structure. By aggregating these local uncertainties, one can determine the global uncertainty of the predicted structure. The inclusion of uncertainties in these predictions offers insights into the confidence and reliability of the model’s outputs. Finally, DeepRelax employs a numerical Euclidean distance geometry (EDG) solver to determine the relaxed structure that satisfies the relaxation quantities. For more details, please refer to Section 4.
@@ 2.2 Benchmark on X-Mn-O dataset
For our initial benchmarking, we utilize the X-Mn-O dataset, a hypothetical elemental substitution database previously employed for photoanode application studiesnoh2019unveiling ; kim2020generative . This dataset derives from the MP database, featuring prototype ternary structures that undergo elemental substitution with X elements (Mg, Ca, Ba, and Sr). It consists of 28,579 data pairs, with each comprising an unrelaxed structure and its corresponding relaxed state. The dataset is divided into training ($N=22,863$), validation ($N=2,857$), and test ($N=2,859$) sets, adhering to an 8:1:1 ratio. As illustrated in Supplementary Materials (SM), Figure S1, there are significant structural differences between the unrelaxed and relaxed structures within this dataset.
@@
To evaluate model performance, we use the MAE of Cartesian coordinates, bond lengths, lattice parameters, and cell volume to measure the consistency between predicted and ground-truth relaxed structures. Additionally, we calculate the match rate—a measure of how closely predicted relaxed structures align with their ground truth counterparts within a defined tolerance, as determined by Pymatgenong2013python . Detailed descriptions of these metrics are provided in Subsection 4.10.
Table 1 presents the comparative results, showing that DeepRelax greatly outperforms other baselines. Notably, DeepRelax shows a remarkable improvement in prediction accuracy over the Dummy model, with enhancements of 63.06%, 68.30%, 71.49%, 89.63%, and 30.71% across coordinates, bond lengths, lattice, cell volumes, and match rate, respectively. Moreover, DeepRelax surpasses the previous SOTA model, Cryslator, by 8.66% in coordinate prediction, and 45.16% in cell volume estimation. Figure 2 (a) shows the distribution of MAE for coordinates, lattice parameters, and cell volumes as predicted by the Dummy model and DeepRelax. DeepRelax demonstrates a notable leftward skewness in its distribution, signifying a tendency to predict structures that closely approach the relaxed state. To visualize the performance of Deeprelax, we take two typical structures, $\rm Sr_{4}Mn_{2}O_{6}$ and $\rm Ba_{1}Mn_{4}O_{8}$ from the X-Mn-O database (see Figure 3), and relax them using DeepRelax. As can be seen, the predicted structures are highly consistent with the DFT-relaxed ones. The results demonstrate close agreement with DFT-relaxed structures. More randomly selected samples are validated by DFT calculations in Subsection 2.5.
Table 1: Comparative results of DeepRelax and baseline models on the X-Mn-O dataset, based on MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), cell volume (Å$\rm\AA^{3}$), and match rate (%) between predicted and ground truth relaxed structures
@@
Figure 2: Distribution of MAE for predicted structures by the Dummy model and DeepRelax across datasets: (a) X-Mn-O, (b) MP, and (c) C2DB. The subfigures display the MAE for coordinates, lattice parameters, and cell volumes, respectively.
Figure 3: Visualization of two crystal structures relaxed by DeepRelax: (a) $\rm Sr_{4}Mn_{2}O_{6}$ and (b) $\rm Ba_{1}Mn_{4}O_{8}$, where a, b, c, $\alpha$, $\beta$, and $\gamma$ are lattice constants. The results demonstrate close agreement with DFT-relaxed structures.
@@ 2.3 Benchmark on Materials Project
To demonstrate DeepRelax’s universal applicability across various elements of the periodic table and diverse crystal types, we conduct further evaluations using the Materials Project datasetchen2022universal . This dataset spans 89 elements and comprises 187,687 snapshots from 62,783 compounds captured during their structural relaxation processes. By excluding compounds missing either initial or final structures, we refined the dataset to 62,724 pairs. Each pair consists of an initial and a corresponding final structure, providing a comprehensive basis for assessing the performance of DeepRelax. This dataset is then split into training, validation, and test data in the ratio of 90%, 5%, and 5%, respectively. As illustrated in Figure S1, the structural differences for each pair tend toward a mean absolute error (MAE) of zero, indicating that many initial structures are closely aligned with their final counterparts.
Training a direct machine learning (ML) model for datasets with varied compositions poses significant challenges, as evidenced in Cryslatorkim2020generative . This model shows reduced prediction performance when trained on the diverse MP database. Despite these challenges, DeepRelax demonstrates its robustness and universality. As indicated in Table 2, DeepRelax significantly surpasses the three baseline models in coordinate prediction, highlighting its effectiveness even in diverse and complex datasets. Figure 2 (b) shows the MAE distribution for predicted structures compared to the ground truth for the MP dataset, which is less significant compared to the results for the X-Mn-O dataset shown in Figure 2 (a). This is because many initial structures closely resemble their final states in the MP database as evidenced by Figure S1. Consequently, the MP dataset presents a more complex learning challenge for structural relaxation models.
Table 2: Comparison of results between the proposed DeepRelax and other models on the MP dataset. The improvement is calculated by comparing DeepRelax with the Dummy model
@@ 2.4 Transfer learning of pre-trained DeepRelax on Computational 2D Materials Database
To demonstrate the universal application of DeepRelax, we extend the application of DeepRelax, initially pre-trained on 3D materials from the MP dataset, to 2D materials through transfer learning. We take the C2DB datasethaastrup2018computational ; gjerding2021recent ; lyngby2022data as an example, which covers 62 elements and comprises 11,581 pairs of 2D crystal structures, each consisting of an initial and a final structure. The dataset is divided into training, validation, and testing subsets, maintaining a ratio of 6:2:2. The structural differences for each pair in this dataset fall within the range observed for the X-Mn-O and MP datasets, as shown in Figure S1.
In this context, DeepRelax trained via transfer learning is denoted as DeepRelaxT to differentiate it from DeepRelax. Table 3 illustrates our key findings: Firstly, both DeepRelax and DeepRelaxT outperform the other three baselines in the C2DB dataset, proving the applicability of our direct ML model to the system of 2D materials. Figure 2 (c) presents the MAE distribution for predicted structures by the Dummy model and DeepRelax on the C2DB dataset. These results suggest a modest improvement over the Dummy model for the C2DB dataset when compared to other datasets: the improvement is more significant than that observed for the MP dataset as depicted in Figure 2 (b). Secondly, DeepRelaxT demonstrates notable improvements over DeepRelax, with enhancements of 5.61% in coordinates, 38.43% in bond length, 3.53% in lattice, and 5.81% in cell volume in terms of MAE. Finally, DeepRelaxT shows a faster convergence rate than DeepRelax, as detailed in SM Section 2, Figure S2. These results underline the benefits of large-scale pretraining and the efficacy of transfer learning.
Table 3: Training DeepRelax on the C2DB dataset via transfer learning

### v2

@@ Abstract
In computational molecular and materials science, determining equilibrium structures is the crucial first step for accurate subsequent property calculations. However, the recent discovery of millions of new crystals and complex twisted structures has challenged traditional computational methods, both ab initio and machine-learning-based, due to their computationally intensive iterative processes. To address these scalability issues, here we introduce DeepRelax, a deep generative model capable of performing geometric crystal structure relaxation rapidly and without iterations. DeepRelax learns the equilibrium structural distribution, enabling it to predict relaxed structures directly from their unrelaxed ones. The ability to perform structural relaxation at the millisecond level per structure, combined with the scalability of parallel processing, makes DeepRelax particularly useful for large-scale virtual screening. We demonstrate DeepRelax’s reliability and robustness by applying it to five diverse databases, including oxides, Materials Project, two-dimensional materials, van der Waals crystals, and crystals with point defects. DeepRelax consistently shows high accuracy and efficiency, validated by density functional theory calculations. Finally, we enhance its trustworthiness by integrating uncertainty quantification. This work significantly accelerates computational workflows, offering a robust and trustworthy machine-learning method for material discovery and advancing the application of AI for science. Code for DeepRelax is available at https://github.com/Shen-Group/DeepRelax.
@@ 1 Introduction
Atomic structural relaxation is usually the first step and foundation for further computational analysis of properties in computational chemistry, physics, materials science, and medicine. This includes applications such as chemical reactions on surfaces, complex defects in semiconductor heterostructures, and drug design. To date, computational relaxation algorithms have typically been achieved using iterative optimization, such as traditional ab initio methods as shown in Fig. 1(a). For example, each iteration in density functional theory (DFT) calculations involves solving the Schrödinger equation to determine the electronic density distribution, from which the total energy of the system can be calculated. The forces on each atom, derived from differentiating this energy with respect to atomic positions, guide atomic movements to lower the system’s energy, typically using optimization algorithms. Despite its effectiveness, the high computational demands and poor scalability of DFT limit its applications across high-dimensional chemical and structural spaces zuo2021accelerating , such as the complex chemical reaction surfaces, doped semiconductor interfaces, or in the structural relaxation of the 2.2 million new crystals recently identified by DeepMind merchant2023scaling . It is worth noting that the discovery of huge new materials has been significantly accelerated by high-throughput DFT calculations ong2013python ; saal2013materials ; curtarolo2012aflow ; zhou20192dmatpedia and advanced machine learning (ML) algorithms chen2021phase ; xie2021crystal ; zeni2023mattergen ; zhao2023physics , which is promoting the development of more efficient relaxation algorithms.
@@
ML has emerged as a promising alternative for predicting relaxed structures chen2022universal ; deng2023chgnet ; zuo2021accelerating ; mosquera2024machine ; kolluru2022open ; kim2023structure ; yoon2020differentiable ; wang2024concurrent ; Omee2024 . As conventional iterative optimization, iterative ML approaches chen2022universal ; deng2023chgnet ; zuo2021accelerating ; kolluru2022open ; mosquera2024machine ; wang2024concurrent ; Omee2024 utilize surrogate ML models to approximate energy and forces at each iteration, as shown in Fig. 1(a), thereby circumventing the need to solve the computationally intensive Schrödinger equation. A typical example is the defect engineering in crystalline materials kazeev2023sparse ; mosquera2023identifying ; huang2023unveiling . Mosquera-Lois et al. mosquera2024machine and Jiang et al. jiang2024machine demonstrated that ML surrogate models could accelerate the optimization of crystals with defects. These ML models can retain DFT-level accuracy by training on extensive databases containing detailed information on structural relaxations, including energy, forces, and stress.
@@
However, there are two primary challenges in current iterative ML structural optimizers: training data limitations and non-scalability. Their training dataset must include full or partial intermediate steps of DFT relaxation. However, almost all publicized material databases, such as ICSD belsky2002new and 2DMatPedia zhou20192dmatpedia , do not provide such structural information, potentially limiting the application of iterative ML methods. The other challenge is that the large-scale parallel processing capability of iterative ML methods is limited due to their step-by-step nature. To address this, Yoon et al. yoon2020differentiable developed a model called DOGSS, and Kim et al. kim2023structure proposed a model named Cryslator. Both conceptually introduce direct ML approaches to predict the final relaxed structures from their unrelaxed counterparts. However, these approaches have only been validated on specific datasets or systems, and their universal applicability to diverse datasets or systems remains unproven.
@@
Figure 1: An overview of ML methods for crystal structure relaxation. (a) Iterative ML methods that iteratively estimate energy and force to determine the equilibrium structure. (b) Illustration of our proposed DeepRelax method, which employs a periodicity-aware equivariant graph neural network (PaEGNN) to directly predict the relaxation quantities. Euclidean distance geometry (EDG) is then used to determine the final relaxed structure that satisfies the predicted relaxation quantities.
@@
In this work, we introduce DeepRelax, a scalable, universal, and trustworthy deep generative model designed for direct structural relaxation. DeepRelax requires only the initial crystal structures to predict equilibrium structures in just a few hundred milliseconds on a single GPU. Furthermore, DeepRelax can efficiently handle multiple crystal structures in parallel by organizing them into mini-batches for simultaneous processing. This capability is especially advantageous in large-scale virtual screening, where rapid assessment of numerous unknown crystal configurations is essential. To demonstrate the reliability and robustness, we evaluate DeepRelax across five different datasets, including diverse 3D and 2D materials: the Materials Project (MP) jain2013commentary , X-Mn-O oxides kim2020generative ; kim2023structure , the Computational 2D Materials Database (C2DB) haastrup2018computational ; gjerding2021recent ; lyngby2022data , layered van der Waals crystals, and 2D structures with point defects huang2023unveiling ; kazeev2023sparse . DeepRelax not only demonstrates superior performance compared to other direct ML methods but also exhibits competitive accuracy to the leading iterative ML model, M3GNet chen2022universal , while being approximately 100 times faster in terms of speed. Moreover, we conduct DFT calculations to assess the energy of DeepRelax’s predicted structures, confirming our model’s ability to identify energetically favorable configurations. Additionally, DeepRelax employs an uncertainty quantification method to assess the trustworthiness of the model. Finally, we would like to highlight that the aim of using DeepRelax is not to replace DFT relaxation, but to make the predicted structures close to the DFT-relaxed configuration. Thus, the DFT method can rapidly complete the relaxation with only a few steps, significantly speeding up the traditional ab initio relaxation process, especially for complex structures.
@@ 2.1 DeepRelax architecture
DeepRelax emerges as a solution to the computational bottlenecks faced in DFT methods for crystal structure relaxation. Fig. 1(b) shows the workflow of DeepRelax, which takes an unrelaxed crystal structure as input and uses a periodicity-aware equivariant graph neural network (PaEGNN) to predict the relaxation quantities, including interatomic distances in the relaxed structure, displacements between the initial and relaxed structures, and the lattice matrix of the relaxed structure. DeepRelax then employs a numerical Euclidean distance geometry (EDG) solver to determine the relaxed structure that satisfies the predicted relaxation quantities. In addition, DeepRelax also quantifies bond-level uncertainty for each predicted interatomic distance and displacement. Aggregating these bond-level uncertainties allows for the computation of the system-level uncertainty, offering valuable insights into the trustworthiness of the model.
A notable feature of PaEGNN, distinguishing it from previous graph neural networks (GNNs) schutt2021equivariant ; xie2018crystal , is the explicit differentiation of atoms in various translated cells to encode periodic boundary conditions (PBCs) using a unit cell offset encoding (UCOE). Additionally, its design ensures equivariance, facilitating active exploration of crystal symmetries and thus providing a richer geometric representation of crystal structures.
@@ 2.2 Benchmark on X-Mn-O dataset
For our initial benchmarking, we utilize the X-Mn-O dataset, a hypothetical elemental substitution database previously employed for photoanode application studies noh2019unveiling ; kim2020generative . This dataset derives from the MP database, featuring prototype ternary structures that undergo elemental substitution with X elements (Mg, Ca, Ba, and Sr). It consists of 28,579 data pairs, with each comprising an unrelaxed structure and its corresponding DFT-relaxed state. The dataset is divided into training ($N=22,863$), validation ($N=2,857$), and test ($N=2,859$) sets, adhering to an 8:1:1 ratio. As illustrated in Suppl. Fig. 1, there are significant structural differences between the unrelaxed and DFT-relaxed structures within this dataset.
@@
To evaluate model performance, we use the mean absolute error (MAE) of Cartesian coordinates, bond lengths, lattice matrix, and cell volume to measure the consistency between predicted and DFT-relaxed structures. Additionally, we calculate the match rate—a measure of how closely predicted relaxed structures align with their ground truth counterparts within a defined tolerance, as determined by Pymatgen ong2013python . Detailed descriptions of these metrics are provided in Subsection 4.10.
Table 1 presents the comparative results, showing that DeepRelax greatly outperforms other baselines. Notably, DeepRelax shows a remarkable improvement in prediction accuracy over the Dummy model, with enhancements of 63.06%, 68.30%, 71.49%, 89.63%, and 30.71% across coordinates, bond lengths, lattice, cell volumes, and match rate, respectively. Moreover, DeepRelax surpasses the previous SOTA model, Cryslator, by 8.66% in coordinate prediction, and 45.16% in cell volume estimation. Fig. 2(a) shows the distribution of MAE for coordinates, lattice matrix, and cell volumes as predicted by the Dummy model and DeepRelax. DeepRelax demonstrates a notable leftward skewness in its distribution, signifying a tendency to predict structures that closely approach the DFT-relaxed state. To visualize the performance of DeepRelax, we take two typical structures, $\rm Sr_{4}Mn_{2}O_{6}$ and $\rm Ba_{1}Mn_{4}O_{8}$ from the X-Mn-O database (see Fig. 3), and relax them using DeepRelax. As can be seen, the DeepRelax-predicted structures are highly consistent with the DFT-relaxed ones. The results demonstrate close agreement with DFT-relaxed structures. More DFT validations are in Subsection 2.7.
Table 1: Comparative results of DeepRelax and baseline models on the X-Mn-O dataset, evaluated based on MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), cell volume (Å$\rm\AA^{3}$), and match rate (%) between the predicted and DFT-relaxed structures
@@
The best performance in each metric is highlighted in bold.
@@
Figure 2: Distribution of MAE for predicted structures by the Dummy model and DeepRelax. (a) X-Mn-O dataset, (b) MP dataset, and (c) C2DB dataset. Each subfigure, from left to right, displays the MAE for coordinates ($\rm\AA$), lattice matrices ($\rm\AA$), and cell volumes (Å$\rm\AA^{3}$), respectively. Source data are provided with this paper.
Figure 3: Visualization of two crystal structures relaxed by DeepRelax. (a) $\rm Sr_{4}Mn_{2}O_{6}$ and (b) $\rm Ba_{1}Mn_{4}O_{8}$, where $a$, $b$, and $c$ are lattice constants in angstroms (Å), and $\alpha$, $\beta$, and $\gamma$ are angles in degrees (). The results demonstrate close agreement between DeepRelax-predicted structures and DFT-relaxed structures.
@@ 2.3 Benchmark on Materials Project
To demonstrate DeepRelax’s universal applicability across various elements of the periodic table and diverse crystal types, we conduct further evaluations using the Materials Project dataset chen2022universal . This dataset spans 89 elements and comprises 187,687 snapshots from 62,783 compounds captured during their structural relaxation processes. By excluding compounds missing either initial or DFT-relaxed structures, we refined the dataset to 62,724 pairs. Each pair consists of an initial and a corresponding DFT-relaxed structure, providing a comprehensive basis for assessing the performance of DeepRelax. This dataset is then split into training, validation, and test data in the ratio of 90%, 5%, and 5%, respectively. As illustrated in Suppl. Fig. 1, the structural differences for each pair tend toward an MAE of zero, indicating that many initial structures are closely aligned with their DFT-relaxed counterparts.
Training a direct ML model for datasets with varied compositions poses significant challenges, as evidenced in Cryslator kim2023structure . This model shows reduced prediction performance when trained on the diverse MP database. Despite these challenges, DeepRelax demonstrates its robustness and universality. As indicated in Table 2, DeepRelax significantly surpasses the three baseline models in coordinate prediction, highlighting its effectiveness even in diverse and complex datasets. Fig. 2(b) shows the MAE distribution for predicted structures compared to the DFT-relaxed ones for the MP dataset, which is less significant compared to the results for the X-Mn-O dataset shown in Fig. 2(a). This is because many initial structures closely resemble their DFT-relaxed structures in the MP database as evidenced by Suppl. Fig. 1. Consequently, the MP dataset presents a more complex learning challenge for structural relaxation models.
Table 2: Comparison of results between the proposed DeepRelax and other models on the MP dataset. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures. The improvement is calculated by comparing DeepRelax with the Dummy model
@@
The best performance in each metric is highlighted in bold.
@@ 2.4 Transfer learning of pre-trained DeepRelax on Computational 2D Materials Database
To demonstrate the reliable application of DeepRelax, we extend the application of DeepRelax, initially pre-trained on 3D materials from the MP dataset, to 2D materials through transfer learning. We take the C2DB dataset haastrup2018computational ; gjerding2021recent ; lyngby2022data as an example, which covers 62 elements and comprises 11,581 pairs of 2D crystal structures, each consisting of an initial and a DFT-relaxed structure. The dataset is divided into training, validation, and testing subsets, maintaining a ratio of 6:2:2. The structural differences for each pair in this dataset fall within the range observed for the X-Mn-O and MP datasets, as shown in Suppl. Fig. 1.
In this application, DeepRelax trained via transfer learning is denoted as DeepRelaxT to differentiate it from DeepRelax. Table 3 illustrates our key findings: Firstly, both DeepRelax and DeepRelaxT outperform the other three baselines in the C2DB dataset, proving the applicability of our direct ML model to 2D materials. Fig. 2(c) presents the MAE distribution for predicted structures by the Dummy model and DeepRelax on the C2DB dataset. These results suggest a modest improvement over the Dummy model. Notably, this improvement surpasses those observed for the MP dataset as depicted in Fig. 2(b). Secondly, DeepRelaxT demonstrates notable improvements over DeepRelax, with enhancements of 5.61% in coordinates, 38.43% in bond length, 3.53% in lattice, and 5.81% in cell volume in terms of MAE. Finally, DeepRelaxT shows a faster convergence rate than DeepRelax, as detailed in Suppl. Fig. 2. These results underline the benefits of large-scale pretraining and the efficacy of transfer learning.
Table 3: Comparison of results among DeepRelax, DeepRelaxT (transfer learning version), and other models on the C2DB dataset. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures
@@
The best performance in each metric is highlighted in bold.
2.5 Application in layered vdW crystals
Layered vdW crystals are of significant interest in the field of materials science and nanotechnology because of their unique tunable structures, such as twisting and sliding configurations li2022deep . One notable characteristic of these crystals is that the weak inter-layer vdW force may significantly change upon full relaxation, while the strong intra-layer chemical bonds undergo relatively small changes.
To demonstrate the reliable performance of our DeepRelax model on this type of crystal, we performed DFT relaxation of 58 layered vdW crystals covering 29 elements using van der Waals corrections, parameterized within the DFT-D3 Grimme method. Given the small sample size, we employ transfer learning, utilizing a model pre-trained on the Materials Project dataset.
Suppl. Table 1 shows the inter-layer distances for the unrelaxed, DFT-D3-relaxed, and DeepRelax-predicted structures of six vdW layered crystals. The inter-layer distances of the predicted structures closely match those of the relaxed structures, highlighting the effectiveness of transferred DeepRelax on layered vdW crystals. Furthermore, an analysis of the MAE in bond length for representative bonding pairs, detailed in Suppl. Table 2, further demonstrates DeepRelax’s precision in predicting structural changes in layered vdW crystals.
2.6 Application in crystals with defects
Most crystals have intrinsic defects. To demonstrate the robustness of DeepRelax to crystal structures with neutral point defects, we employ $\rm MoS_{2}$ structures with a low defect concentration, including 5,933 different defect configurations within an 8 $\times$ 8 supercell, as cataloged by Huang et al. huang2023unveiling , to evaluate DeepRelax. Suppl. Fig. 3 demonstrates a notably lower MAE in both atom coordinates and bond lengths for DeepRelax compared to the Dummy model, thereby underscoring DeepRelax’s robustness and efficacy in defect structure calculations, which is further validated by DFT calculations in the next chapter.
Figure 4: DFT validations. (a) Distributions of deviations for 100 random samples from the X-Mn-O dataset, measured using MAE in coordinates ($\rm\AA$) between the unrelaxed and DFT-relaxed structures. (b) Energy distribution for the three types of structures among the 100 random samples from the X-Mn-O dataset. The boxplots show the median (black line inside the box), interquartile range (box), and whiskers extending to 1.5 times the interquartile range, with outliers plotted as individual points. (c) Distribution of deviations for 100 random samples from the Materials Project dataset with relatively rational initial structures. (d) Energy distribution for the three types of structures among the 100 random samples from the Materials Project dataset. (e) Distributions of deviations for 20 random samples from the 2D materials defect dataset. (f) The number of DFT ionic steps required, starting from the initial unrelaxed structures and the DeepRelax-predicted structures, respectively. The samples are sorted based on the number of ionic steps required by the unrelaxed structures for better observation. Source data are provided with this paper.